### v1

# organization=Université de Rennes, IRISA – UMR 6074,addressline=263 Av. Général Leclerc, city=Rennes, postcode=35000, country=France

Abstract
In this article, we give an extended space formulation for the induced tree and path polytopes of chordal graphs with variables associated with the edge and vertex sets. Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities. We show which of these inequalities define facets and exhibit a superset of the facet-defining ones that can be enumerated in polynomial time. We show that for some graphs, the latter superset contains redundant inequalities. As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs .
Keywords— induced path , induced trees , extended formulations , polytope , chordal graphs
Definitions
All the graphs in this article are simple and connected unless specified. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We respectively denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, let $E(W)$ denote the set of edges of $G$ having both endpoints in $W$ and $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an extremity are said to be adjacent. For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. A vertex is simplicial if its neighborhood induces a clique.
An induced path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ). A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). The length of a path or hole is its number of edges. A path or hole is called long if its length is at least 4. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior. A tree is a graph with no hole.
The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
A graph is $k$-chordal if its holes have maximum size $k$, 3-chordal graphs are called chordal. Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number. A tree is a connected graph containing no holes.
Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by dim $\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is included in the $n$-dimensional hypercube and all its extreme points are binary. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$.
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by the extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V}\times\{0,1\}^{E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0. Given a point = $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)_{G^{\prime}}$ its restriction to elements of $G^{\prime}$. A cone is a polyhedron described by a zero right-handed linear system; it is pointed if it contains no affine space. The elements of a cone can be described by conic combinations of finite sets of elements called generating. The minimal generating families of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators.
Motivations
In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral results associated with these problems.
Induced trees
The problem of finding the size of the largest induced tree in a graph is a problem that has been approached through extremal combinatorics [ERDOS198661, MATOUSEK2007307, FOX2009494]. In particular, an important question is that of bounding this number according to several graph parameters, such as the number of vertices, edges, radius, independence number, maximum clique, and connectivity. In [ERDOS198661], a Ramsey-type bound is proposed on the minimum integer $n(k,t)$ such that every graph having at least $n(k,t)$ vertices contains either a clique of size $k$ or an induced tree of size $t$. Finally, they show that every connected graph with radius $r$ has an induced tree of size at least 2r12r-1 by exhibiting a large induced path. A related problem is the one of finding large induced forests [inducedforest] in which bounds for the largest induced forest are given with respect to maximum degree or stability number. Besides these studies, no polyhedral study of the induced tree polytope on particular classes has been proposed. The problem of computing the maximum size of an induced tree is known to be NP-hard [completePathTree] in general, but it is also the case in bipartite and triangle-free graphs with maximal degree 3 [MATOUSEK2007307]. Several integer linear programs have been proposed to find maximum weighted such subsets [ILPinducedtree], and extensive computational experiments have been made on general graphs.
Induced paths
The problem of finding a longest induced path has applications in assessing the robustness of a network. It can be viewed as a variant of the graph diameter. In a graph where vertices represent communication devices that may fail, and edges represent possible transmission links, the length of the longest induced path indicates the worst-case number of intermediate devices required for a one-to-one communication between any two vertices $u$ and $v$ after the failure of a set of devices that do not disconnect $u$ from $v$. Another application is about error correcting codes [error]. The induced paths of a graph are induced trees, and the type of questions raised in the previous paragraph can be raised considering this structure. The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of size at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of given size $k$. In [esperet], the authors show that every 3-connected planar graph contains an induced path on $\text{log}(|V|)$ vertices. In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of size $n$ to force the existence of an induced path of size $log(n)$. This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $(log n)^{\Omega(1/tlog^{2}t)}$.
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees. Finally, extended space ILPs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
Contributions
In this paper, we first exhibit an extended formulation of compact size characterizing the induced tree polytope of chordal graphs, showing that the problem of computing a maximum vertex and/or edge weighted induced tree is solvable in polynomial time in chordal graphs. For a given graph, the induced tree polytope contains the induced path polytope, and both characterizations follow similar schemes of proof.
Even if the maximum vertex weighted induced path is already known to be solvable in polynomial time in chordal graphs [GAVRIL2002203, HANSEN2009135] it is unknown whether it is the case for vertex and edge weighted graphs. For this reason, we provide an extended formulation for the induced path polytope. The set of inequalities of this formulation is exponential, and we characterize which of them define facets. Finally, we show that a polynomial superset of the facet-defining ones can be enumerated efficiently. This exhibits a compact formulation for the induced path polytope of chordal graphs. Note that both latter results are necessary as our characterization of the facet-defining ones does not show that they come in a polynomial number.
1 Induced tree polytope of chordal graphs
In this section, we give an extended formulation for the induced tree polytope of a chordal graph. Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{T}(G)$ the following system of inequalities:
$\sum_{u\in K}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
$-x_{v}$ $\leq 0$ $\forall v\in V$ (2)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (3)
$x(V)-y(E)$ $=1$ (4)
We now prove that this system fully characterizes induced tree polytopes in chordal graphs.
Theorem 1.1.
A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
Proof. $\Rightarrow"$ Note that the system of inequalities (1)–(3) defines a pointed cone $\mathcal{C}(G)$. The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}(G)$’s generators that satisfy (4). Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}(G)$ and satisfy (4). We show that they are the generators of $\mathcal{C}(G)$ by induction on a simplicial decomposition of $G$.
To initialize the induction, consider the graph $G$ with a single vertex, $\mathcal{C}(G)$ is the positive half line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that $\mathcal{C}(H)$ is generated by induced trees’ extended incidence vectors for $H=G\setminus\{v\}$. Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $\mathcal{C}(H)$ by construction. By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced trees $\Gamma$ with coefficients $\lambda$. The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$. Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}$ but increments either $y_{uv}^{\prime}$ or $x_{v}^{\prime}$. We first modify the conic combination so that $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in E$. We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$. Let us consider the following subsets of $\Gamma$:
  • $\Gamma^{u}\ \forall u\in N(v)$ containing its elements whose intersection with $N[v]$ equals $\{u\}$.
  • $\Gamma^{u}_{2}\ \forall u\in N(v)$ containing its elements whose intersection with $N(v)$ is $\{u,w\}$ for some $w\in N(v)$.
  • $\Gamma^{u}_{v}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$
We refer to Figure 1 for an illustration of the sets of induced trees considered. Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
To each element $T$ of $\Gamma^{u}$, $T\cup\{v\}$ is an induced tree of $G$. This makes it possible to apply the Algorithm (1) over each element of $\bigcup_{u\in N(v)}\Gamma^{u}$, incrementing $y_{uv}^{\prime}$ and $x_{u}^{\prime}$.
Let $T$ be an element of $\Gamma^{u}$
$\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$
$T^{\prime}=T\cup\{v\}$
$\lambda_{T}=\lambda_{T}-\alpha$
$\lambda_{T^{\prime}}=\lambda_{T^{\prime}}+\alpha$
$T$ is removed from $\Gamma$ if $\alpha=\lambda_{T}$
$T^{\prime}$ is added to $\Gamma$
Algorithm 1 Reduction of $\Gamma^{u}$
Algorithm 1 increases $y_{uv}^{\prime}$ unless $y_{uv}^{\prime}=y_{uv}^{*}$ or $\Gamma^{u}$ is empty.
Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}(G)$ we have the following:
$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{u}_{v})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction. This show how $y_{uv}^{*}=y_{uv}^{\prime}$ can be achieved on all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ and exhibit a conic comination of $(x^{*},y^{*})$.
$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $C$ induce a hole of size at least 4 in $G$, and $v$ belongs to $C$. By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(C)$ and $x_{v}=2$ and 0 on all other components, we obtain a point which belongs to $\mathcal{T}(G)$ and is not included in the hypercube implying that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary. ∎
[…]
Figure 1: Sets of induced trees considered in the proof of Theorem 1.1
Corollary 1.2.
The induced tree polytope of a graph $G$ in the extended variable space is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal.
The set of inclusion-wise maximal cliques of chordal graphs is polynomial (1)–(3), implying that Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
Corollary 1.3.
The edge and vertex weighted maximum induced tree is solvable in polynomial time on chordal graphs.
2 Induced path polytope of chordal graphs
We first present a technical lemma concerning edge contraction in chordal graphs exhibited in [dupontbouillard2024contractions]. After that, we give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated such constraints come in a polynomial number.
Lemma 2.1 ([dupontbouillard2024contractions]).
If $G=(V,E)$ is chordal, then $G/F$ also is chordal for any set $F\subseteq E$.
Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$, we simply write $C(W)$ when clear from context. We denote by $\mathcal{P}(G)$ the following system:
$y(\delta(w,C(K\cup\{w\})))+2y(\delta(w,K))$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
$x(V)-y(E)$ $=1$ (8)
The next theorem shows that $\mathcal{P}(G)$ describes the induced path polytope of chordal graphs. Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$) and the extremitie of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient one. At most two green edges belong to a same induced path every blue edge is not contained in any induced path containing some other colored edge.
Theorem 2.2.
A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary:
Proof. $\Rightarrow$ We proceed by showing that the pointed cone $\mathcal{C}(G):=$(5)–(7) is generated by the extended incidence vectors of induced paths. The extreme points of (5)–(8) are scalings of the generators of $\mathcal{C}(G)$ satisfying (8).
We now show that the set of induced paths’ extended induced path generates $\mathcal{C}(G)$. To do so, we proceed by induction on a perfect elimination order of $G$. Initialization of the induction is given by the single vertex graph $G$ for which $\mathcal{C}(G)=\mathds{R}_{+}$.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose $\mathcal{C}(H)$ to be generated by $H$’s induced path extended incidence vectors for $H=G\setminus v$. Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $C(H)$.
By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced path $\Psi$ with coefficients $\lambda$. The proof consists of iteratively modifying $\Psi$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Psi$ and $\lambda$; an induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0. Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}=(x^{*},y^{*})_{H}$. Let us fix an edge $uv\in\delta(v)$, we exhibit a conic combination whose $y_{uv}$ component is equal to $y_{uv}^{*}$, as applying the procedure will preserve the value of $y_{e}$ for the other edge $e\in\delta(v)\setminus\{uv\}$, it can be applied until the conic combination is equal to $(x^{*},y^{*})$ on all components but $x_{v}$.
Let us consider the following subsets of $\Psi$:
  • $\Psi^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and do not contain any other element of $N[v]$
  • $\Psi_{2}^{u}$ is the subset of $\Psi$ having $u$ as an extremity and containing another element of $N(v)$
  • $\Psi^{u}_{I}$ is the subset of $\Psi$ whose interior contains $u$ and does not contain $v$
  • $\Psi_{u}^{v}$ is the subset of $\Psi$ containing both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
  • $\Psi_{\bot}$ contains the remaining elements.
We refer to Figure 3 for an illustration of the different types of path considered. To each element of $\Psi^{u}$, the addition of $v$ yields another induced path which contains $\{u,v\}$. This implies that unless $y_{uv}^{\prime}$ is already equal to $y_{uv}^{*}$, a conic combination preserving $(x^{\prime},y^{\prime})_{H}$ and strictly increasing over $y_{uv}$ can be computed. Algorithm 2 describes such an operation.
let $P$ be an element of $\Psi^{u}$
$\alpha=\min\{y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P}\}$
$P^{\prime}=P\cup\{v\}$
$(x^{\prime},y^{\prime})=(x^{\prime},y^{\prime})-\alpha\chi^{P}+\alpha\chi^{P^{\prime}}$
$\lambda_{P}=\lambda_{P}-\alpha$
$\lambda_{P^{\prime}}=\lambda_{P^{\prime}}+\alpha$
Algorithm 2 Reduction of $\Psi^{u}\ \forall u\in N(v)$
We can now suppose that $\Psi^{u}$ is empty (otherwise equality of the conic combination’s $y_{uv}$ component and $y_{uv}^{*}$ is already achieved) which implies the following:
$x^{*}_{u}$ $=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})$ (9)
$y_{uv}^{\prime}$ $=\lambda(\Psi_{u}^{v})$ (10)
Claim 2.3.
If $\Psi_{I}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
Proof. Suppose that $\Psi^{u}_{I}$ is empty. Inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ applied to $(x^{*},y^{*})$ yields: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$. Using equality (9) and (10) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}.$
Suppose now that $\Psi_{I}^{u}$ is non-empty.
let $P$ be an element of $\Psi_{I}^{u}$
let $P_{1}$ and $P_{2}$ be the two different maximal subpath of $P$ ending in $u$.
$\alpha=\min(\lambda_{P},\lambda_{u,v}$)
$\lambda_{P}=\lambda_{P}-\alpha$
Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
$j=i+1\mod 2$
$\lambda_{u,v}=\lambda_{u,v}-\alpha$
$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
Algorithm 3 Removal of $\{u,v\}$ from $\Psi_{u}^{v}$
Algorithm (3) is used to ensure that each element of $\Psi_{u}^{v}$ intersects $\delta(u)$ twice, such a path always contains $uv$ and another edge in $\delta(u)$. When Algorithm (3) cannot be applied, either $\Psi_{I}^{u}$ is empty and Claim (2.3) concludes or $\lambda_{u,v}=0$.
Claim 2.4.
If $\Psi_{2}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
Proof. Suppose that $\Psi_{2}^{u}$ is empty. Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$. Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$. Equalities (9) and (10) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and conclude the proof. ∎
Suppose now that $\Psi_{2}^{u}$ is nonempty, let us consider the clique $K\subseteq N(u)\cap N(v)$ of vertices contained in at least one element of $\Psi_{2}^{u}$. We now consider the partition of $\Psi_{I}^{u}$ into $\Psi_{I_{1}}^{u}$ containing the elements intersecting $K$ and $\Psi_{I_{0}}^{u}$ containing the others. Algorithm 4 permits to remove elements from $\Psi_{I_{1}}^{u}$ and adds elements in $\Psi_{u}^{v}$ or $\Psi_{2}^{u}$.
Let $P$ be in $\Psi_{I_{1}}^{u}$
Decompose $P$ into $P_{1}$ and $P_{2}$ the two maximal subpath of $P$ ending in $u$
$\alpha=\min(y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P})$
$\lambda_{P}=\lambda_{P}-\alpha$
Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
$j=i+1\mod 2$
$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
Algorithm 4 Reduction of $\Psi_{I_{1}}^{u}$
Using Algorithm 4 permits obtaining a conic combination of induced path such that $y_{uv}^{\prime}=y_{uv}^{*}$, or no element of $\Psi_{I}^{u}$ intersects $K$ that is $\Psi_{I}^{u}=\Psi_{I_{0}}^{u}$.
Claim 2.5.
Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$. Either $u_{1}$ is adjacent to $u_{2}$ or $P_{2}\cup P_{1}$ induces a path.
Proof. By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path. Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either they intersect in some other vertex than $u$ or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ by construction. By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$. ∎
[…]
Figure 2: Structure of inequalities (5) for a given vertex $w$ and clique $K$ of its neighborhood.
Claim (2.5) now permits to decompose the elements of $\Psi_{I_{0}}^{u}$ for which a neighbor of $u$ is not complete to $K$ into their two maximal subpaths ending in $v$. We then combine one of these subpaths with an element of $\Psi_{2}^{u}$ to obtain other induced paths.
Algorithm 5 gives the procedure. Note that Algorithm 5 may add elements to $\Psi^{u}$ and $\Psi_{I_{1}}^{u}$ which can be empty again using Algorithm 2 and 4. Note that the fact that Algorithm 2 and 4 strictly increase the value of $y_{uv}^{\prime}$ ensures that the sequence of updates on $\lambda$ and $\Psi$ converges.
Let $P_{0}\in\Psi_{I_{0}}^{u}$ and $P_{2}\in\Psi_{2}^{u}$ be such that the $P_{0}^{\prime}\cup P_{2}$ induces a path where $P_{0}^{\prime}$ is a maximal subpath of $P_{0}$ ending in $u$.
$\alpha=\min(\lambda_{P_{2}},\lambda_{P_{0}})$
$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
$\lambda_{P_{0}}=\lambda_{P_{0}}-\alpha$
$\lambda_{P_{2}\cup P_{0}^{\prime}}=\lambda_{P_{2}\cup P_{0}^{\prime}}+\alpha$
$\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}=\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}+\alpha$
Algorithm 5 Replacing the elements of $\Psi_{I_{0}}^{u}$ so that in the remaining ones the neighbors of $u$ are complete to $K$
Now we suppose that the elements of $\Psi^{u}_{I_{0}}$ are complete to $K$ (otherwise, $\Psi^{u}_{2}$ is empty and Claim 2.4 concludes).
Using another time Claim 2.5, we get rid of elements in $\Psi_{u}^{v}$ for which the neighbor of $u$ distinct from $v$ is not complete to $K$ with Algorithm 6. Note that Algorithm 6 increments $\lambda_{uv}$ which is then decreased back using Algorithm 3. It also increments $\Psi_{I_{1}}^{u}$, which is then decreased back using Algorithm 4 if a path in which a neighbor of $u$ is not complete to $K$ is added. Such a sequence of modifications strictly increases $y_{uv}^{\prime}$, ensuring convergence. Note that the neighbor of $u$ in $P_{i}$ as defined in Algorithm 3 is now complete to $K$ as we already used Algorithm 5.
Let $P_{4}$ be in $\Psi_{u}^{v}\setminus\{u,v\}$ and $P_{2}\in\Psi_{2}^{u}$ such that $P_{4}\cup P_{2}\setminus\{v\}$ induces a path
$\alpha=\min(\lambda_{P_{4}},\lambda_{P_{2}})$
$\lambda_{P_{4}}=\lambda_{P_{4}}-\alpha$
$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
$\lambda_{P_{4}\cup P_{2}\setminus\{v\}}=\lambda_{P_{4}\cup P_{2}\setminus\{v\}}+\alpha$
$\lambda_{u,v}=\lambda_{uv}+\alpha$.
Algorithm 6 The neighbors of $u$ in $G[P]$ for $P\in\Psi^{u}$ are complete to $K$
As $(x^{*},y^{*})$ is a point of $\mathcal{C}(G)$, it satisfies the constraint (5) associated with $K$ and $u$ that is:
$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K))+y^{*}(\delta(u,C(K\cup\{u\})))$ (11)
By construction, each element $P_{4}$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K$ implying that $\lambda_{P_{4}}$ contributes one time in $y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]$. By construction, $y^{*}(\delta(u,C(K\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I}^{u})$ . Using equalities (9) and (10) we get:
$2x_{uv}^{*}-2x_{uv}^{*}\geq 2\lambda(\Psi_{2}^{u})+y^{*}_{uv}+2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})-2\lambda(\Psi_{2}^{u})-2\lambda(\Psi_{I}^{u})-2\lambda(\Psi_{u}^{v})$
Which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$. One can repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$. Setting $\lambda_{\{v\}}=(x^{*}_{u}-y^{*}(\delta(v)))$ yields the desired conic combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
$(\Leftarrow)$ The proof is by contradiction. The same non binary point exhibited in the proof of Theorem 1.1 also belongs to $\mathcal{P}(G)$. ∎
[…]
Figure 3: Induced paths considered in the proof of Theorem 2.2
Note that inequalities (6) are mandatory only when the graph has isolated vertices. If the input graph has no such vertex, then inequalities (5) and (7) ensure that $x$ is positive. These constraints have been added to give easy initializations of the induction. The proof of Theorem 2.2 shows that the extreme points of $\mathcal{P}(G)$ are induced paths’ extended incidence vectors, which implies the following.
Corollary 2.6.
A graph $G$ is chordal if and only if its extended space induced path polytope is equal to $\mathcal{P}(G)$.
It turns out that not all inequalities (5) define facets, we characterize which do.
Theorem 2.7.
Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\in N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K)]$ is bipartite.
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K)]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K,\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 8 concludes:
  1. 1.
    $\xi^{v}\ \forall v\in V\setminus\{w\}$
  2. 2.
    $\xi^{u,v}\ \forall uv\in E\setminus\delta(w)$
  3. 3.
    $\xi^{u,w}\ \forall uw\in\delta(w,K)$
  4. 4.
    $\xi^{u,w,v}\ \forall uw\in\delta(w)\setminus\delta(K)\setminus\delta(C(K))$ for some vertex $v\in K$ non adjacent to $u$
The remaining $|\delta(w,C(K))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K)]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$. This can be done by starting from an odd cycle and then adding the missing edges for this set to be spanning the vertices. Now, match each edge of $\delta(w,C(K))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K)|=|\delta(w,C(K))|$. Let $uw$ be in $\delta(w,C(K))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non adjacent hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. This construction is necessary as such path will contain two edges of $\delta(w,C(K))$ making it hard to build an indentity of wanted size over these components.
We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and ”?” stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof. The rank of $H$ is unkown, however, we know that the restriction of the last $|\delta(w,C(K))|$ points to elements of $G\setminus\delta(w,C(K))$ contains a incidence matrix of a non bipartite graph $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K)|$.
Linear independence is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the latter set of points forms a submatrix of rank $|C(K)|$ which are the only ones involving elements of $\delta(w,C(K))$. ∎
[…]
Figure 4: Matrix of points built in the proof of Theorem 2.7 where $G^{\prime}$ is the graph with vertex set $V$ and edge set $(E\setminus\delta(w))\cup(\delta(w)\cap\delta(K))$
The characterization of Theorem 2.7 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently. Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, if $K\subseteq K^{\prime}$ then we have $C(K^{\prime})\subseteq C(K)$, for this reason, we say that a clique defines an orbit if there exists no $K^{\prime}$ strictly containing $K$ such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
Theorem 2.8.
Given a chordal graph $G$, the set of non-dominated inequalities inequalities (5) has polynomial size.
Proof. We build a chordal auxiliary graph $H$ to $G[N(w)]$, whose maximal cliques are in bijection with orbit-defining cliques of $G[N(w)]$. Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will conclude.
The auxiliary graph $H$ is built from $G[N(w)]$ by first adding a true twin to each vertex of $N(w)$, let $W$ be the set of added twins. Secondly, add all the edges between elements of $W$. A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$. This proves that $H$ is chordal.
Let $K_{H}$ be a maximal clique of $H$. It is partitioned into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Now, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{k}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. ∎
Corollary 2.9.
The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
It turns out that the target of Theorems 2.7 and 2.8 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.7 implies that its associated inequalities are not facet defining.
11223344556677
Figure 5: An orbit defining clique whose associated inequalities (5) are not facet defining
Conclusion
In this paper, we propose extended formulations for the induced tree and path polytopes of chordal graphs of compact size, implying the polynomial solvability of finding a maximum vertex and edge weighted such induced subgraph in chordal graphs. Our proofs rely heavily on the structure of the induced trees and paths cones; it may be interesting to investigate the structures of the linear programs whose constraint sets correspond to incidence vectors of induced paths and trees.
Acknowledgment
The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support.