### v2

# Extended formulations for induced tree and path polytopes of chordal graphs

Abstract
In this article, we give two extended space formulations, respectively, for the induced tree and path polytopes of chordal graphs with vertex and edge variables. These formulations are obtained by proving that the induced tree and path extended incidence vectors of chordal graphs form Hilbert basis. This also shows that both polytopes have the integer decomposition property in chordal graphs.
Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities. We show which of these inequalities define facets and exhibit a superset of the facet-defining ones that can be enumerated in polynomial time. We show that for some graphs, the latter superset contains redundant inequalities. As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs.
keywords:
[…]
Definitions
All the graphs in this article are simple and connected unless specified. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We respectively denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, let $E(W)$ denote the set of edges of $G$ having both endpoints in $W$ and $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an endpoint are said to be adjacent. Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in both $W_{1}$ and $W_{2}$. For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when clear from context. A vertex is simplicial if its neighborhood induces a clique.
An induced path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ). Note that sets of vertices inducing cliques with three vertices are considered holes. An induced tree is a set of vertices inducing an acyclic graph. A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). The length of a path or hole is its number of edges. A path or hole is called long if its length is at least 4. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior. A tree is a graph with no hole.
The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$; parallel edges and loops are removed. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
A graph is $k$-chordal if its holes have maximum size $k$; 3-chordal graphs are called chordal. Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number.
The size of a linear system is its number of rows. Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by $\textrm{dim}~\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is bounded and all its extreme points are binary. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$. A cone is a polyhedron described by a linear system with a right hand side equal to 0; it is pointed if it contains no affine space. The elements of a cone can be described by conic combinations of finite sets of elements called generators. The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators they are in one to one correspondence with its face of dimension 1. A polyhedron $\mathcal{P}$ is said to have the integer decomposition property (IDP) if every integer point in the dilation $k\mathcal{P}$ for $k\in\mathds{Z}_{+}$ is obtained as the sum of $k$ elements of $\mathcal{P}$. A set of points $X=\{\chi_{1},\dots,\chi_{n}\}$ forms a Hilbert basis if every integer point in the conic hull of $X$ is expressible as an integer nonnegative combination of elements of $X$. The Hilbert basis and integer decomposition property are closely related. If the extreme points of a polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has the IDP.
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by its extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V\cup E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0. Given a point $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
In this paper, given a graph $G$, we consider the following polyhedra:
  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
Motivations
In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral literature associated with these problems.
Induced trees
The problem of finding the size of the largest induced tree in a graph is a problem that has been approached through extremal combinatorics [ERDOS198661, MATOUSEK2007307, FOX2009494]. In particular, an important question is that of bounding this number according to several graph parameters, such as the number of vertices, edges, radius, independence number, maximum clique, and connectivity. In [ERDOS198661], a Ramsey-type bound is proposed on the minimum integer $n(k,t)$ such that every graph having at least $n(k,t)$ vertices contains either a clique of size $k$ or an induced tree of size $t$. Finally, they show that every connected graph with radius $r$ has an induced tree of size at least 2r12r-1 by exhibiting a large induced path. A related problem is the one of finding large induced forests [inducedforest] in which bounds for the largest induced forest are given with respect to maximum degree or stability number. Besides these studies, no polyhedral study of the induced tree polytope on particular classes has been proposed. The problem of computing the maximum size of an induced tree is known to be NP-hard [completePathTree] in general, but it is also the case in bipartite and triangle-free graphs with maximal degree 3 [MATOUSEK2007307]. Several integer linear programs have been proposed to find maximum weighted such subsets [ILPinducedtree], and extensive computational experiments have been made on general graphs.
Induced paths
The problem of finding a longest induced path has applications in assessing the robustness of a network. It can be viewed as a variant of the graph diameter. In a graph where vertices represent communication devices that may fail, and edges represent possible transmission links, the length of the longest induced path indicates the worst-case number of intermediate devices required for a one-to-one communication between any two vertices $u$ and $v$ after the failure of a set of devices that do not disconnect $u$ from $v$. Another application is about error correcting codes [error]. The induced paths of a graph are induced trees, and the type of questions raised in the previous paragraph can be raised considering this structure. The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of size at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of given size $k$. In [esperet], the authors show that every 3-connected planar graph contains an induced path on $\text{log}(|V|)$ vertices. In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of size $n$ to force the existence of an induced path of size $log(n)$. This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $(logn)^{\Omega(1/tlog^{2}t)}$.
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees. Finally, extended space integer linear programs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
Contributions
In this paper, we study both the induced tree polytope and the induced path polytope of chordal graphs in two separate sections. It turns out that beside being included one in the other, these two polytopes have similar properties. Each section starts by respectively showing that the induced path and tree extended incidence vectors of chordal graphs form a Hilbert basis of some cones, which we explicitly describe by linear systems of inequalities. We show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space. While the system we provide for the induced tree polytope is easily seen to be compact in chordal graphs, the one we provide for the induced path polytope has an exponential number of inequalities. Therefore, we characterize the ones that define facets and show that a superset can be enumerated in polynomial time. This implies that there exists a compact system describing the induced path polytope of chordal graphs. Finally, these results imply that the problems of finding a maximum vertex and edge weight induced tree or path are polynomially solvable in chordal graphs.
1 Induced tree polytope of chordal graphs
In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis. Then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs. Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
$-x_{v}$ $\leq 0$ $\forall v\in V$ (2)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (3)
Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}_{T}(G)$.
Theorem 1.1.
The induced tree extended incidence vectors of a chordal graph form a Hilbert basis.
Proof. We show that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ form a Hilbert basis of $\mathcal{C}_{T}(G)$. As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each one-dimensional face contains integer points. This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$.
We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that the integer points of $\mathcal{C}_{T}(G\setminus v)$ are obtained as positive integer combinations of induced trees’ extended incidence vectors. Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, $(x^{*},y^{*})|_{G\setminus v}$ is an integer point of $\mathcal{C}_{T}(G\setminus v)$ by construction. By the induction hypothesis, let $(x^{*},y^{*})|_{G\setminus v}$ be generated by the induced trees $\Gamma$ with integer coefficients $\lambda$. The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$. For fixed $uv\in\delta(v)$, each operation preserves the value of $(x^{\prime},y^{\prime})|_{H}$ for $H=G\setminus uv$, and preserves nonnegativity and integrality of $\lambda$. We define the slackness $\Delta=\sum_{uv\in\delta(v)}\Delta(u,v)$ for $\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}$. We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$.
Let us consider the following subsets of $\Gamma$, illustrated in Figure 1 :
  • $\Gamma^{u}\ \forall u\in N(v)$ containing its elements whose intersection with $N[v]$ equals $\{u\}$.
  • $\Gamma^{u}_{2}\ \forall u\in N(v)$ containing its elements whose intersection with $N(v)$ is $\{u,w\}$ for some $w\in N(v)$.
  • $\Gamma^{u}_{v}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$
Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
We define the following operation to be applied on $(\Gamma,\lambda)$, which decreases $\Delta$ by an integer value.
$\mathbf{\Gamma^{u}}$-Reduction: let $T$ be an element of $\Gamma^{u}$. Set $\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$ increase $\lambda_{T\cup\{v\}}$ and decrease $\lambda_{T}$ by $\alpha$.
By construction, for any $T\in\Gamma^{u}$, $T\cup\{v\}$ induces a tree and that $\mathbf{\Gamma^{u}}$-Reduction preserves integrality of $\lambda$. Moreover, $\xi^{T}|_{H}=\xi^{T\cup\{v\}}|_{H}$ implying that $\lambda(\Psi)|_{H}$ is preserved.
$\mathbf{\Gamma^{u}}$-Reduction applies until $\Gamma^{u}$ is empty or $\Delta(u,v)=0$.
Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{u}_{v})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction. This show how $\Delta(u,v)=0$ can be achieved on all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with integer $\lambda$.
Note that Theorem 1.1 is not a characterization of the graphs for which the induced tree’s extended incidence vectors form a Hilbert basis. In fact, this statement is also true for $C_{4}$. However, it is straightforward to see that this property is closed by taking induced subgraphs. A natural question would be to characterize the graphs for which this property holds.
Theorem 1.1 implies the following.
Corollary 1.2.
The induced tree polytope and cone of a chordal graphs have IDP.
[…]
Figure 1: Sets of induced trees considered in the proof of Theorem 1.3
Let us denote by $\mathcal{T}(G)$ the polytope defined as the intersection of $\mathcal{C}_{T}(G)$ and the following hyperplane:
$\{(x,y):x(V)-y(E)=1\}$ (4)
We now prove that $\mathcal{T}(G)$ characterizes induced tree polytopes in chordal graphs.
Theorem 1.3.
A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
Proof. $\Rightarrow"$ The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators that belong to (4). Theorem 1.1 concludes.
$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $C$ induce a hole of size at least 4 in $G$, and $v$ belongs to $C$. By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(C)$ and $x_{v}=2$ and 0 on all other components, we obtain a point which belongs to $\mathcal{T}(G)$ and is not included in the hypercube implying that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary. ∎
The set of inclusion-wise maximal cliques of chordal graphs is polynomial, implying that (1)–(3) has compact size. Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
Corollary 1.4.
The induced tree polytope of a graph $G$ is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal. Moreover, $\mathcal{T}(G)$ is described by a linear system of size $O(|V||\mathcal{K}|)$.
Corollary 1.5.
The edge and vertex weighted maximum induced tree is solvable in polynomial time on chordal graphs.
2 Induced path polytope of chordal graphs
We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions]. We use this result to prove a technical graph theory lemma.
After that, we show that the extended incidence vectors of induced paths of a chordal graph form a Hilbert basis. This allows us to give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated constraints come in a polynomial number.
Lemma 2.1 ([dupontbouillard2024contractions]).
If $G=(V,E)$ is chordal, then $G/F$ is also chordal for any set $F\subseteq E$.
The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis is constructive; therefore, it is necessary to build induced paths with "good properties," and this is what the following lemma stands for.
Claim 2.2.
Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$. Either $u_{1}$ is adjacent to $u_{2}$ or $P_{2}\cup P_{1}$ induces a path.
Proof. By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path. Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either they intersect in some other vertex than $u$ or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths. By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$.
To show that a set of vectors forms a Hilbert basis, it can be easier to know the polyhedral description of the cone they generate. Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system and show in the next theorem that it is generated by the Hilbert basis formed by the induced paths’ extended incidence vectors.
[…]
Figure 2: Structure of inequalities (5) for a given vertex $w$ and clique $K$ of its neighborhood.
$y(\delta(w,C(K\cup\{w\})))+2y(\delta(w,K))$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient 1. At most two green edges belong to a same induced path; every blue edge is not contained in any induced path containing some other colored edge. This shows that extended incidence vectors of induced paths belong to $\mathcal{C}_{P}(G)$.
2.1 Induced paths of chordal graphs and Hilbert basis
In this section, $G$ denotes a chordal graph, $v$ a simplicial vertex, $u$ a neighbor of $v$ and $H$ the graph obtained by removing $uv$ from $G$. We denote by $\Psi$ a family of vertex subsets of $G$ inducing paths, and consider $\lambda$ to be an integer vector indexed by the induced paths of $G$. The pair $(\Psi,\lambda)$ acts as the pair $(\Gamma,\lambda)$ used in the previous section.
We prove that the extended incidence vectors of induced paths of chordal graphs form a Hilbert basis. Although the proof follows the same inductive scheme as in the case of induced trees, it is substantially more technical. In particular, it relies on a refined decomposition of the family $\Psi$.
We first introduce this decomposition of $\Psi$, then define five operations acting on $(\Psi,\lambda)$. For each operation, we establish a corresponding lemma showing that it preserves the property that all elements of $\Psi$ induce paths of $G$, the nonnegativity of $\lambda$ and the value of $\lambda(\Psi)|_{H}:=\sum_{P\in\Psi}\lambda_{P}\xi^{P}|_{H}$. Moreover, we precise its impact on the coefficients $\lambda$. Finally, these ingredients are combined to complete the proof.
Let us consider the following partition of $\Psi$:
  • $\Psi^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and do not contain any other element of $N[v]$
  • $\Psi_{2}^{u}$ is the subset of $\Psi$ having $u$ as an extremity and containing another element of $N(v)$
  • $\Psi^{u}_{I}$ is the subset of $\Psi$ whose interior contains $u$ and does not contain $v$
  • $\Psi_{u}^{v}$ is the subset of $\Psi$ containing both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
  • $\Psi_{\bot}$ contains the remaining elements.
[…]
Figure 3: Types of induced paths considered in Section 2.1
We refer to Figure 3 for an illustration of the different types of path considered, it may be used as a visual aid to keep track of the definitions . We denote by $K_{\Psi}$ the elements of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$:
$K_{\Psi}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}$
$K_{\Psi}$ is considered to be changing accordingly to $\Psi$. We further partition $\Psi_{I}^{u}$ into:
  • $\Psi_{K_{\Psi}}^{u}$ containing the elements of $\Psi$ intersecting $K_{\Psi}$
  • $\Psi_{\overline{K}_{\Psi}}^{u}$ containing the others.
We now define some operations to be applied on ($\Psi$$\lambda)$. We say that such an operation is valid if it preserves the integrality and nonnegativity of $\lambda$, if the vertex subsets added to $\Psi$ induce paths, and if it preserves the value of $\lambda(\Psi)|_{H}$. In each reduction, $\alpha$ is chosen as an integer satisfying $0<\alpha\leq\min\{\lambda_{P}:P\in\Psi,\ \lambda_{P}\text{ is decreased}\}$. This ensures that $\lambda$ stays nonnegative.
$\mathbf{\Psi^{u}}$-Reduction : let $P$ be an element of $\Psi^{u}$. Increase $\lambda_{P\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
$\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{K_{\Psi}}^{u}$. $P$ contains the two inclusion-wise maximal distinct subsets $P_{1}$ and $P_{2}$ inducing paths ending in $u$. Let $P_{2}$ be the one intersecting $N(v)\setminus u$. Increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
$\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$ such that a neighbor $u_{1}$ of $u$ in $P$ is not complete to $K_{\Psi}$. Let $P_{1}$ be its inclusion-wise maximal subset inducing a path ending in $u$ and containing $u_{1}$, and let $P^{\prime}$ be the other inclusion-wise maximal subset inducing a path ending in $u$. Let $P_{2}$ be an element of $\Psi_{2}^{u}$ whose element $u_{2}$ in $K_{\Psi}$ is non-adjacent to $u_{1}$. Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ by $\alpha$.
$\mathbf{\Psi_{u}^{v}}$-Reduction : Let $P$ be a path of $\Psi_{u}^{v}$ such that the neighbor $u_{1}$ of $u$ distinct from $v$ is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$. Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$. Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$.
$\mathbf{\lambda_{\{u,v\}}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$. $P$ contains the two distinct inclusion-wise maximal subset $P_{1}$ and $P_{2}$ both inducing paths ending in $u$. Note that $P_{1}\cup\{v\}$ induces a path. Decrease $\lambda_{u,v}$ and increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$.
Note that in each of the above operations, $\alpha$ is integer, implying that they preserve integrality of $\lambda$. We show that under some conditions, the above reductions are valid. It remains to verify that the sets of vertices built induce paths and that the value of $\lambda(\Psi)|_{H}$ is preserved.
Lemma 2.3.
Suppose that $\Psi^{u}$ is nonempty. $\mathbf{\Psi^{u}}$-Reduction is valid.
Proof. Given $\in\Psi^{u}, PP$ ends in $u$ and does not intersect $N(v)\setminus\{u\}$, therefore $P\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P\cup\{v\}}|_{H}$. ∎
Lemma 2.4.
Suppose that $\Psi_{K_{\Psi}}^{u}$ is non-empty. $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and strictly increments $\lambda(\Psi^{u}_{2})$, $\lambda(\Psi^{v}_{u})$.
Proof. By construction, $P\in\Psi_{K_{\Psi}}^{u}$ is decomposed into its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ that induce paths ending in $u$. At most one of them intersects $N(v)\setminus\{u\}$, which implies that $P_{1}\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}.$ Therefore, $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and increments both $\lambda_{P_{1}\cup\{v\}}$ and $\lambda_{P_{2}}$. ∎
Lemma 2.5.
Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ contains an element $P$ such that a neighbor $u_{1}\in P$ of $u$ is not complete to $K_{\Psi}$, in particular $\Psi_{2}^{u}$ is non empty. Then $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid.
Proof. Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$, let $P^{\prime}$ be its other subset inducing a path ending in $u$. Let $P_{2}$ be an element of $\Psi^{u}_{2}$ such that the neighbor $u_{2}$ of $u$ in $P_{2}$ is non adjacent to $u_{1}$. Claim 2.2 implies that $P_{1}\cup P_{2}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}\cup P_{2}}|_{H}+\xi^{P^{\prime}\cup\{v\}}|_{H}.$ Therefore $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid. ∎
Lemma 2.6.
Suppose that $\Psi_{u}^{v}$ contains an element $P$ such that the neighbor $u_{1}$ of $u$, distinct from $v$, is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$. $\mathbf{\Psi_{u}^{v}}$-Reduction is valid and strictly increases $\lambda_{\{u,v\}}$.
Proof. Let $P_{1}=P\setminus\{v\}$, by Claim 2.2 $P_{1}\cup P_{2}$ induces a path and $\{u,v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}}|_{H}+\xi^{\{u,v\}}|_{H}$. ∎
Lemma 2.7.
Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non empty. $\mathbf{\lambda_{\{u,v\}}}$-Reduction is valid and strictly decreases $\lambda_{\{u,v\}}$.
Proof. Given $P\in\Psi_{\overline{K}_{\Psi}}^{u}$, we consider its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ inducing a path ending in $u$. $P_{1}\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{\{u,v\}}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}$. ∎
The proof of the following theorem proceeds by induction over the number of vertices and makes use of the previous definitions and operations.
Theorem 2.8.
The extended incidence vectors of induced paths of a chordal graph form a Hilbert basis of $\mathcal{C}_{P}(G)$.
Proof. The statement is trivial for $|V|=1$. Induction goes as follows: As $G$ is chordal, it contains a simplicial vertex $v$. The graph $G\setminus v$ is chordal and by hypothesis, the induced paths’ extended incidence vectors of $G\setminus v$ form a Hilbert basis of $\mathcal{C}_{P}(G\setminus v)$. Let $(x^{*},y^{*})$ belong to $\mathcal{C}_{P}(G)\cap\mathds{Z}^{V\cup E}$, its restriction $(x^{*},y^{*})|_{G\setminus v}$ belongs to $\mathcal{C}_{P}(G\setminus v)\cap\mathds{Z}^{V\cup E}$. From now on, $(\Psi,\lambda)$ denotes an arbitrary family of induced paths of $G$ with integer coefficients satisfying $(x’,y’)|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$, where (x,y)(x’,y’) is the point induced by $(\Psi,\lambda)$. All operations defined above transform $(\Psi,\lambda)$ into another such family and preserve: integrality of $\lambda$, the fact that the elements of $\Psi$ induce paths of $G$ and $(x’,y’)|_{H}=(x^{*},y^{*})|_{H}$. Note that the latter requirement implies that the operations not only preserve the value of $\lambda(\Psi)$ over elements in $G\setminus v$ but also on edges in $\delta(v)\setminus\{uv\}$.
The goal of the upcoming sequence of reductions is to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$. For this we consider the slackness;
$\Delta:=\sum_{uv\in\delta(v)}\Delta(u,v)\text{ where }\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}.$
The proof proceeds by fixing a neighbor $u$ of $v$ and modifying $(\Psi,\lambda)$ until the equality $\Delta(u,v)=0$ holds. Since reductions for a fixed $u$ do not affect $\Psi_{w}^{v}$ for $w\neq u$, and that elements of $\Psi^{u}_{w}$ are the only ones contributing to the value of $y_{vw}^{\prime}$, the procedure can be applied independently for each neighbor of $v$.
An induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0.
Lemmas 2.32.5 give conditions for some valid reductions strictly decreasing $\Delta(u,v)$ to be applied for respectively $\alpha=\min(\Delta(u,v),\lambda_{P})$, $\alpha=\min(\Delta(u,v),\lambda_{P})$, and $\alpha=\min(\lambda_{P},\lambda_{P_{2}},\Delta(u,v))$. Since the corresponding reductions strictly decrease either $\Delta(u,v)$, at some point, none of Lemma 2.32.6 apply for the fixed neighbor $u$. This implies that, either $\alpha=0$ and $\Delta(u,v)=0$ which ends the proof or the following holds:
  • $\Psi^{u}$, $\Psi^{u}_{I_{1}}$ both are empty
  • the neighbors of $u$ in the elements of $\Psi^{u}_{I_{0}}$ are complete to $K_{\Psi}$.
By construction, the only elements of $\Psi$ which contribute to the value of $y_{uv}^{\prime}$ belong to $\Psi_{u}^{v}$, therefore the following holds :
$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
As $\Psi^{u}$ is empty, the following holds:
$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{\overline{K}_{\Psi}}^{u})+\lambda(\Psi_{u}^{v})\\$ (9)
Suppose that $\Psi^{u}_{I_{0}}$ is empty. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$. Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta(u,v)=0$.
Assume henceforth that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non-empty.
Suppose $\Psi_{2}^{u}$ to be empty, Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$. Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$. Equalities (9) and (8) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and $\Delta(u,v)=0$.
We may now assume that $\Psi_{2}^{u}$ is nonempty.
Therefore, $K$ is non-empty and the neighbors of $u$ in the elements of $\Psi_{\overline{K}_{\Psi}}^{u}$ are complete to $K_{\Psi}$.
Suppose that the conditions of Lemma 2.6 hold, then applying $\Psi_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$ yields an element of $\Psi_{K_{\Psi}}^{u}$, therefore, the conditions of Lemma 2.4 hold and $\Psi_{K_{\Psi}}^{u}$-Reduction applies. This sequence of operations strictly decreases $\Delta(u,v)$.
Suppose that the conditions of Lemma 2.7 hold, applying $\lambda_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{u,v}$) yields an element of $\Psi_{u}^{v}$, and one of $\Psi^{u}$. Therefore, the conditions of Lemma 2.3 hold, and $\Psi^{u}$-Reduction applies. This sequence of operations also decreases $\Delta(u,v)$.
As $\Psi^{u}$-Reduction, $\Psi_{K_{\Psi}}^{u}$-Reduction, $\Psi_{\overline{K}_{\Psi}}^{u}$-Reduction, and the two latter sequences of reductions strictly decrease $\Delta(u,v)$, we may assume that none of Lemmas 2.32.7 apply or that $\Delta(u,v)=0$. Therefore, the neighbors of $u$ in the elements of $\Psi^{v}_{u}$ are complete to $K_{\Psi}$ and $\lambda_{\{u,v\}}=0$.
As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is (where $C(S)$ denotes the set of vertices complete to $S$):
$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K_{\Psi}))+y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))$ (10)
By construction, each element $P$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K_{\Psi}$. This implies that $\lambda_{P}$ contributes two times in $y^{*}[\delta(u,C(K_{\Psi}\cup\{u\}))]$. By construction, $y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K_{\Psi}\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{\overline{K}_{\Psi}}^{u})$. Using equalities (8) and (9) we get:
$0\geq y^{*}_{uv}-\lambda(\Psi_{u}^{v})$
which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$. Repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$. Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired integer combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
As induced trees of $C_{4}$ coincide with its induced path, Theorem 2.8 is not a characterization of the graph for which the induced path extended incidence vectors have IDP.
2.2 Extended formulation for the induced path polytope
In this section, we discuss the consequences of Theorem 2.8 and use it to give a compact description of the induced path polytope of chordal graphs.
Corollary 2.9.
The induced path cone and polytope of chordal graphs both have IDP.
Similarly to what we do for the induced tree polytope, we exhibit a description of the induced path polytope of chordal graphs. We denote by $\mathcal{P}(G)$ the polytope defined as the intersection of $\mathcal{C}_{P}(G)$ and the hyperplane:
$\{(x,y)~|~x(V)-y(E)=1\}.$ (11)
Theorem 2.10.
A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary.
Proof. $(\Rightarrow)$ The extreme points of $\mathcal{P}(G)$ are scalings of $\mathcal{C}_{P}(G)$’s generators which belong to (11). Theorem 2.8 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11), which concludes the proof.
$(\Leftarrow)$ The proof is by contradiction. The same non binary point exhibited in the proof of Theorem 1.3 also belongs to $\mathcal{P}(G)$. ∎
Note that inequalities (6) are mandatory only when the graph has isolated vertices. If the input graph has no such vertex, then inequalities (5) and (7) ensure that $x$ is positive. The proof of Theorem 2.10 shows that the extreme points of $\mathcal{P}(G)$ are induced paths’ extended incidence vectors, which implies the following.
Corollary 2.11.
A graph $G$ is chordal if and only if its induced path polytope is equal to $\mathcal{P}(G)$.
It turns out that not all inequalities (5) define facets; we characterize which ones do.
Theorem 2.12.
Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\subseteq N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K\cup\{w\})]$ is bipartite.
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K\cup\{w\},\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 11 concludes. We consider $K_{w}=K\cup\{w\}$. The points are as follows:
  1. 1.
    $\xi^{\{v\}}\ \forall v\in V\setminus\{w\}$
  2. 2.
    $\xi^{\{u,v\}}\ \forall uv\in E\setminus\delta(w)$
  3. 3.
    $\xi^{\{u,w\}}\ \forall uw\in\delta(w,K)$
  4. 4.
    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus(\delta(K)\cup\delta(C(K_{w})))$ for some vertex $v\in K$ non adjacent to $u$
The remaining $|\delta(w,C(K_{w}))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$. This can be done by starting from an odd cycle and then adding the missing edges for this set to span the vertices. Now, match each edge of $\delta(w,C(K_{w}))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K_{w})|=|\delta(w,C(K_{w}))|$. Let $uw$ be in $\delta(w,C(K_{w}))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non-adjacent, hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. This construction is necessary as such a path will contain two edges of $\delta(w,C(K_{w}))$, making it impossible to continue a diagonal equal to 1. We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and "?" stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof.
$H$ is similar to the incidence matrix of the non bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
Full row rank is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the last set of points forms a submatrix of rank $|C(K_{w})|$ which are the only ones involving elements of $\delta(w,C(K_{w}))$. This implies that the matrix has rank $|V|+|E|$, therefore the inequality is facet defining. ∎
[…]
Figure 4: Matrix of points built in the proof of Theorem 2.12 where $G^{\prime}$ is the graph with vertex set $V$ and edge set $(E\setminus\delta(w))\cup(\delta(w)\cap\delta(K))$
The characterization of Theorem 2.12 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently. Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, let $K_{w}=K\cup\{w\}$, $K_{w}^{\prime}=K^{\prime}\cup\{w\}$. If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$, for this reason, we say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
Theorem 2.13.
Given a chordal graph $G$, the set of non-dominated inequalities (5) has polynomial size.
Proof. We build a chordal auxiliary graph $H$ to $G[N(w)]$, whose maximal cliques are in bijection with orbit-defining cliques of $G[N(w)]$. Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will conclude.
The auxiliary graph $H$ is built from $G[N(w)]$ by first adding a true twin to each vertex of $N(w)$, let $W$ be the set of added twins. Secondly, add all the edges between elements of $W$. A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$. This proves that $H$ is chordal.
Let $K_{H}$ be a maximal clique of $H$. It is decomposed into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Now, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{K}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. ∎
Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$. Summing over all $w\in V$, the total number of non-dominated inequalities (5) is $O(|V|^{2})$.
Corollary 2.14.
There exists an extended formulation for the induced path polytope of chordal graphs with size $O(|V|^{2})$. The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
It turns out that the target of Theorems 2.12 and 2.13 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.12 implies that its associated inequalities are not facet defining.
11223344556677
Figure 5: An orbit defining clique whose associated inequalities (5) are not facet defining
Conclusion
In this paper, we show that the induced tree and path extended incidence vectors of chordal graphs form a Hilbert basis. We use this fact to give a compact linear system of inequalities whose binaricity characterizes chordality. It turns out that there exist non-chordal graphs for which induced tree and path extended incidence vectors form a Hilbert basis, which raises the question of characterizing such graphs. In general, showing that a set of vectors $\mathcal{V}$ forms a Hilbert basis is used to show that a linear system whose rows are elements of $\mathcal{V}$ is TDI. In this paper, we are not considering the programs whose rows are extended induced path or induced tree of chordal graphs. Finding some problems that are naturally formulated as integer linear programs whose linear constraint set has such a structure would lead to TDI-ness results.
Statements and Declarations
The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support and declares that he has no conflicts of interest.