### v3

# Extended formulations for induced tree and path polytopes of chordal graphs

Abstract
In this article, we give an extended space formulation for the induced tree polytope and another one for the induced path polytope of chordal graphs, both with vertex and edge variables. These formulations are obtained by proving that, in chordal graphs, the extreme points of those polytopes respectively form two Hilbert bases. Since these polytopes are binary, it also shows that they have integer decomposition property.
While the formulation for the induced tree polytope has a polynomial size, the formulation we propose for the induced path polytope involves an exponential number of inequalities. We identify which of these inequalities are facet‑defining and exhibit a polynomially enumerable superset containing all of them, although this superset may include redundant inequalities for some graphs.
As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs.
keywords:
[…]
Definitions
All the graphs in this article are simple. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, we denote by $E(W)$ the set of edges of $G$ having both endpoints in $W$, and by $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$. Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in $W_{1}$ and the other in $W_{2}$. For $F\subseteq E$, we denote by $V(F)$ the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. We denote by $G\setminus v$ (resp. $G\setminus uv$¨) the obtained from $G$ by removing the vertex $v$ (resp. edge $uv$). Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when the graph is clear from context. A clique is a set of pairwise adjacent vertices. A stable set is a set of pairwise nonadjacent vertices. A graph whose vertex set can be partitioned into two stable sets is called bipartite. A vertex is simplicial if its neighborhood induces a clique.
A path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ and $p\geq 4$). A cycle is a set of edges inducing a connected graph with vertices of degree equal to 2. A tree is an acyclic graph.
A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). By abuse of definition, we consider that a set of vertices inducing a path is an induced path, similarly, an induced tree is a set of vertices inducing a tree. The length of a path or hole is its number of edges. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior moreover, we say that $P$ ends in $w_{1}$ and $w_{k}$. The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$; parallel edges and loops are removed. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
A graph is $k$-chordal if its holes have maximum size $k$; graphs with no holes are called chordal. Equivalently, a graph is chordal if there exists an order $(v_{1},\dots,v_{k})$ of its vertices such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$; such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $\{N_{G_{i}}[v_{i}]:i\in\{1,\dots,k\}\}$, which shows that their number is linear.
Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by $\textrm{dim}~\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. The size of a linear system of inequalities is its number of rows. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$, it is binary if all its extreme points are binary. A cone is a polyhedron described by a linear system with a right-hand side equal to 0; it is pointed if it contains no affine space. Given a set of points $P\subseteq\mathds{R}^{n}$, we denote by $\text{cone}(P)$ the conic hull of $P$’s elements, that is, the minimal cone containing $P$. A set of points $P$ is a generating set of the cone $\mathcal{C}$ if $\text{cone}(P)=\mathcal{C}$. The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we call the elements of such a set its generators they are in one-to-one correspondence with its faces of dimension 1. A polyhedron $\mathcal{P}$ is said to have the integer decomposition property (IDP) if every integer point in the dilation $k\mathcal{P}$ for $k\in\mathds{Z}_{+}$ is obtained as the sum of $k$ elements of $\mathcal{P}$. A set of points $X=\{\chi_{1},\dots,\chi_{n}\}$ forms a Hilbert basis if every integer point in the conic hull of $X$ is expressible as an integer nonnegative combination of elements of $X$. Hilbert basis and integer decomposition property are closely related. If the extreme points of a binary polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has IDP.
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. An induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with $E^{\prime}=E(V^{\prime})$, can be encoded by its incidence vector $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the remaining components equal 0. In this article, we consider the extended incidence vector denoted by $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V}\times\{0,1\}^{E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the remaining components equal 0. Considering extended descriptions such as proposed here is a common way to obtain compact polyhedral descriptions of combinatorial polytopes [exntededfo]. In particular, considering extended spaces sometimes allows to obtain compact descriptions, whereas the natural variable may require exponentially many constraints. Given a point $(x,y)\in\mathds{R}^{V}\times\mathds{R}^{E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
Given a graph $G=(V,E)$, the polytopes we consider in this article are defined as follows:
  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ is an induced tree of $G$}\}$
  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ is an induced tree of $G$}\}$
  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ is an induced path of $G$}\}$
  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ is an induced path of $G$}\}$
Notations
Here and later, calligraphic letters denote polyhedra. Capital letters denote sets of vertices or edges, or graphs (note that $\delta$ is also used for sets of edges). Lowercase letters refer to vertices, edges or numbers. The symbol $\Gamma$ is reserved for sets of induced trees, while $\Psi$ for sets of induced paths, $\Omega$ for sets of cliques and $\lambda$ refers to vector of integer coefficients. Finally, $\Delta$ refers to gaps which we aim to reduce to 0.
Motivations
In this article, we study the problems of finding maximum weight induced trees and paths where both the vertices and the edges are weighted. In particular, we fill a gap in the polyhedral literature associated with these problems.
Induced trees
The problem of finding the size of the largest induced tree in a graph is a problem that has been approached through extremal combinatorics [ERDOS198661, MATOUSEK2007307, FOX2009494]. In particular, an important question is that of bounding this number according to several graph parameters, such as the number of vertices, edges, radius, independence number, maximum clique, and connectivity. In [ERDOS198661], a Ramsey-type bound is proposed on the minimum integer $n(k,t)$ such that every graph having at least $n(k,t)$ vertices contains either a clique of size $k$ or an induced tree of size $t$. Finally, they show that every connected graph with radius $r$ has an induced tree of size at least 2r12r-1 by exhibiting a large induced path. A related problem is the one of finding large induced forests [inducedforest] in which bounds for the largest induced forest are given with respect to maximum degree or stability number. Besides these studies, no polyhedral study of the induced tree polytope on particular classes has been proposed. The problem of computing the maximum size of an induced tree is known to be NP-hard [completePathTree] in general, but it is also the case in bipartite and triangle-free graphs with maximal degree 3 [MATOUSEK2007307]. Several integer linear programs have been proposed for the maximum weight induced tree problem [ILPinducedtree] based on different variable spaces, and extensive computational experiments have been made on general graphs.
Induced paths
The problem of finding a longest induced path has applications in assessing the robustness of a network. It can be viewed as a variant of the graph diameter. In a graph where vertices represent communication devices that may fail, and edges represent possible transmission links, the length of the longest induced path indicates the worst-case number of intermediate devices required for a one-to-one communication between any two vertices $u$ and $v$ after the failure of a set of devices that do not disconnect $u$ from $v$. Another application is about error correcting codes [error]. The induced paths of a graph are induced trees, and the type of questions raised in the previous paragraph can be raised considering this structure. The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of length at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of size $k$. In [esperet], the authors show that every 3-connected planar graph contains an induced path on $\text{log}(|V|)$ vertices. In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of length $n$ to force the existence of an induced path of length $log(n)$. This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $log(n)^{\Omega(1/tlog^{2}t)}$.
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has however, been considered in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points in $\{0,1\}^{E}$ encode paths between pairs of leaves in a tree. In trees, induced path and path are the same subgraphs, so the latter article describes the induced path (between leaves) polytope of trees with edge variables. Finally, extended space integer linear programs have been proposed for the maximum weight induced path [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
Contributions
In this article, we study both the induced tree polytope and the induced path polytopes of chordal graphs in two separate sections. It turns out that beside being included one in the other, these two polytopes share similar properties. Both following sections follow the same scheme, the first is about induced trees, and the second is about induced path. Each section starts by showing that the extended incidence vectors of their corresponding graph structure form a Hilbert basis in chordal graphs. As a byproduct, we obtain linear systems of inequalities describing the cones they generate in chordal graphs. Finally, we show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space. While the system we provide for the induced tree polytope is easily seen to be compact in chordal graphs, the one we provide for the induced path polytope has an exponential number of inequalities. Therefore, we characterize the ones that define facets and show that a superset can be enumerated in polynomial time. This implies that there exists a compact system describing the induced path polytope of chordal graphs. This study allows to solve the problems of finding a maximum vertex and edge weight induced tree or a maximum weight induced path in polynomial time for chordal graphs.
1 Induced tree polytope of chordal graphs
In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis, and then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs.
1.1 Induced trees of chordal graphs and Hilbert basis
Given a graph $G$, and a vertex $v$ we denote by $\Omega^{v}$ the set of maximal cliques containing $v$ and by $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\Omega^{v}$ (1)
$-x_{v}$ $\leq 0$ $\forall v\in V$ (2)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (3)
For binary points, Inequalities (1) express the fact that no two edges incident to $v$ and contained in a same triangle belong to an induced tree. Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}_{T}(G)$.
Theorem 1.1.
The induced tree extended incidence vectors of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{T}(G)$.
Proof. We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that the integer points of $\mathcal{C}_{T}(G\setminus v)$ are obtained as positive integer combinations of induced trees’ extended incidence vectors. Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, and $(\lambda,\Gamma)$ an integer nonnegative combination of $G$’s induced trees’ extended incidence vectors satisfying the following properties, where $(x^{\prime},y^{\prime})=\sum_{T\in\Gamma}\lambda_{T}\Gamma_{T}$:
  1. (i)
    $(x^{\prime},y^{\prime})\leq(x^{*},y^{*})$
  2. (ii)
    $(x^{\prime},y^{\prime})|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$
  3. (iii)
    $(\lambda,\Gamma)$ minimize the quantity $\begin{array}[]{l}\\ \Delta(\lambda,\Gamma):=\sum_{e\in\delta(v)}\Delta_{e}(\lambda,\Gamma)\\ \text{where }\Delta_{e}(\lambda,\Gamma):=y_{e}^{*}-y_{e}^{\prime}\ \forall e\in\delta(v)\end{array}$ among the combinations satisfying (i)-(ii).
$\Delta_{uv}(\lambda,\Gamma)$ corresponds to a Manhattan distance to $(x^{*},y^{*})$. An integer nonnegative combination $(\lambda,\Gamma)$ satisfying Conditions (i)-(ii) is given by the induction hypothesis applied on the integer point $(x^{*},y^{*})|_{G\setminus v}$ of $\mathcal{C}_{T}(G\setminus v)$ hence, $(\lambda,\Gamma)$ exists.
We first show that $\Delta_{uv}(\lambda,\Gamma)=0\ \forall uv\in E$. Fix $uv\in E$ and $H=G\setminus uv$, let us consider the following disjoints subsets of $\Gamma$:
  • $\Gamma^{u}\ \forall u\in N(v)$ containing its elements whose intersection with $N[v]$ equals $\{u\}$.
  • $\Gamma^{u}_{2}\ \forall u\in N(v)$ containing its elements whose intersection with $N(v)$ is $\{u,w\}$ for some $w\in N(v)$.
  • $\Gamma^{v}_{u}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$.
Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$. We refer to Figure 1 for an illustration of the different types of induced trees considered, it may be used as a visual aid to keep track of the definitions, we provide two different corresponding induced trees in different shades of green; these are the two bottom ones.
Suppose by contradiction that $\Delta_{uv}(\lambda,\Gamma)$ is nonzero and that $\Gamma^{u}$ is nonempty for some neighbor $u$ of $v$. Let $P$ be an element of $\Gamma^{u}$ and set $\alpha=\min(\Delta_{uv}(\lambda,\Gamma),\lambda_{P})$. Note that $P\cup\{v\}$ is an induced tree, therefore decreasing $\lambda_{P}$ and increasing $\lambda_{P\cup\{v\}}$ by $\alpha$ yields a nonnegative integer combination of induced paths’ extended incidence vectors of $G$, say $(\lambda^{\prime},\Gamma^{\prime})$. As $\xi^{P}|_{G\setminus uv}=\xi^{P\cup\{v\}}|_{G\setminus uv}$, $(\lambda^{\prime},\Gamma^{\prime})$ satisfies Conditions (i)-(ii) and $\Delta_{e}(\lambda,\Gamma)=\Delta_{e}(\lambda^{\prime},\Gamma^{\prime})\ \forall e\in\delta(v)\setminus\{uv\}$. Moreover, $\Delta_{uv}(\lambda^{\prime},\Gamma^{\prime})=\Delta_{uv}(\lambda,\Gamma)-\alpha$, Condition (iii) applied to $(\lambda,\Gamma)$ implies that $\alpha=0$. Therefore, either $\Gamma^{u}$ is empty or $\Delta_{uv}(\lambda,\Gamma)=0$. Suppose that $\Gamma^{u}$ is empty, we now show that $\Delta_{uv}(\lambda,\Gamma)=0$ also holds.
Given a family $\Gamma^{\prime}$ of induced trees, we denote by $\lambda(\Gamma^{\prime})$ the sum of $\lambda$’s components associated with elements of $\Gamma^{\prime}$.
As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{v}_{u})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
The first equality is given by the fact that the only elements of $\Gamma$ contributing to $y_{e}^{\prime},$ with $e\in\delta(u,N(v)\setminus\{u\})$ are the ones of $\Gamma^{u}_{2}$ moreover, each element of $\Gamma^{u}_{2}$ contributes to at most one $y_{e}$ for $e\in\delta(u,N(v)\setminus u)$. The inequality is given by inequality (1) associated with vertex $u$ and clique $N[v]$ . The following equality holds by Condition (ii). The two last equalities hold by the facts that the elements of $\Gamma$ contributing to $x_{u}^{\prime}$ are the ones in $\Gamma^{u}\cup\Gamma^{u}_{2}\cup\Gamma^{v}_{u}$, but $\Gamma^{u}$ is supposed to be empty, and that the only elements of $\Gamma$ contributing to $y_{uv}^{\prime}$ are in $\Gamma^{v}_{u}$.
This shows that $\Delta_{uv}(\lambda,\Gamma)=0$ holds for all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with nonnegative integer coefficients $\lambda$ on induced tree’s extended incidence vectors.
Note that Theorem 1.1 is not a characterization of the graphs for which the induced tree’s extended incidence vectors form a Hilbert basis. In fact, this statement is also true for $C_{4}$. Moreover, it is straightforward to see that this property is closed under taking induced subgraphs. A natural question would be to give a forbidden induced subgraph characterization of the graphs for which this property holds. A minimal forbidden such graph is given by the complement graph of a hole of length 7.
1.2 Induced trees of chordal graphs and integral systems
Theorem 1.1 implies the following.
Corollary 1.2.
The induced tree polytope of chordal graphs have IDP.
[…]
Figure 1: Sets of induced trees considered in the proof of Theorem 1.3
Let us denote by $\mathcal{T}(G)$ the polytope defined as the intersection of $\mathcal{C}_{T}(G)$ and the following hyperplane:
$\{(x,y):x(V)-y(E)=1\}$ (4)
Theorem 1.3.
A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
Proof. $\Rightarrow"$ Suppose $G$ to be chordal. The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators which belong to (4). As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each of its one-dimensional face contains integer points. This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$. Therefore, Theorem 1.1 gives that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ generate $\mathcal{C}_{T}(G)$. Extended incidence vectors of induced trees belong to (4) and hence are extreme points of $\mathcal{T}(G)$. It remains to show that $\mathcal{T}(G)$ is bounded. For each vertex $v$ of $G$, there exists a perfect elimination ordering ending in $v$, say $(v_{1},\dots,v_{n},v)$ (it is a consequence of the fact that minimal separators are cliques [Dirac1961]). By definition of perfect elimination orderings, $N(v_{i})\cap\{v_{i+1},\dots,v_{n},v\}$ is a clique and the following inequalities are valid for $\mathcal{C}_{T}(G)$ :
$y(\delta(v_{i},\{v_{i+1},\dots,v_{n},v\}))\leq x_{v_{i}}\quad\forall i\in\{1,\dots,n\}.$
Therefore, summing the previous family of inequalities with x(V)y(E)=1x(V)-y(E)=1 yields $x_{v}\leq 1$. Moreover, $y$ variables are sandwiched between 0 and $x$ variables.
$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $H$ induce a hole of length at least 4 in $G$, and $v$ belongs to $H$. By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(H)$, $x_{v}=2$ and 0 on all other components, we obtain a point $(x,y)$ which belongs to $\mathcal{T}(G)$ and is not included in the hypercube. This implies that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary. ∎
The set of inclusion-wise maximal cliques of chordal graphs being polynomial (1)–(3) has compact size. Therefore, optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
Corollary 1.4.
The induced tree polytope of a graph $G$ is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal. Moreover, $\mathcal{T}(G)$ is described by a linear system of size $O(|V||\mathcal{K}|)$.
Corollary 1.5.
The edge and vertex weighted maximum induced tree is solvable in polynomial time on chordal graphs.
2 Induced path polytope of chordal graphs
We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions] and use this result to prove a technical lemma. Our proof of the fact that induced paths’ extended incidence vectors of chordal graphs form Hilbert bases follows a similar scheme to the one of the previous section.
We define $(\lambda,\Psi)$ in a similar way to what we did with $(\lambda,\Gamma)$ in the proof of Theorem 1.1. However, the proof of the analogous result for induced path requires a closer study of $(\lambda,\Psi)$. Therefore, we first exhibit a set of properties of $(\lambda,\Psi)$ in the form of lemmas before diving into the proof itself.
In the second section, we deduce an exponential size description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated constraints come in a polynomial number.
Lemma 2.1 ([dupontbouillard2024contractions]).
If $G=(V,E)$ is chordal, then $G/F$ is also chordal for any set $F\subseteq E$.
The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis requires to "reorganize" the induced path of a conic combination. We mean by reorganize that we build a new induced path having desirable properties from other induced path. This is what stands for the following lemma.
Claim 2.2.
Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, and such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$. Either $u_{1}$ is adjacent to $u_{2}$ or $P_{2}\cup P_{1}$ induces a path.
Proof. By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path and that $u_{1}$ is not adjacent to $u_{2}$. Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either their intersection contains $u$ and another vertex, or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle with vertex set $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the edges in $E(P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\})$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of length 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths. By chordality of $G$ and Lemma 2.1, $u_{2}$ is adjacent to $u_{1}$, a contradiction.
To show that a set of vectors forms a Hilbert basis, it can be easier to know the polyhedral description of the cone they generate. Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system.
$y\Biggl(\delta\Bigl(w,C(K\cup\{w\})\Bigr)\Biggr)+2y\Biggl(\delta\Bigl(w,K\Bigr)\Biggr)$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
[…]
Figure 2: Structure of inequalities (5) for a given vertex $w$ and clique $K$ of its neighborhood.
For binary points, Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient 1. At most two green edges belong to a same induced path; given a blue edge $e$, for any other colorful edge $e^{\prime}$, there exists a triangle containing both, therefore no induced path $P$ verifies $\{e,e^{\prime}\}\subseteq E(P)$. This shows that extended incidence vectors of induced paths belong to $\mathcal{C}_{P}(G)$.
Observation 2.3.
The induced paths’ extended incidence vectors of a chordal graph $G$ belong to $\mathcal{C}_{P}(G)$.
2.1 Induced paths of chordal graphs and Hilbert basis
This section is dedicated to the proof of the following conjecture.
Conjecture 2.4.
The induced path extended incidence vectors of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{P}(G)$.
We prove Conjecture 2.4 by induction over the vertex set of any chordal graph $G$. Note that the initilization is given by the single vertex graph for which $\mathcal{C}_{P}(G)$ is the positive half line trivially generated by a unit vector. The induction proof consists in considering an integer point $(x^{*},y^{*})$ in $\mathcal{C}_{P}(G)$, a simplicial vertex $v$ of $G$ (which exists by chordality of $G$) and a nonnegative integer combination of $G$’s induced paths’ extended vectors say, the induced path $\Psi$ with coefficient $\lambda$ satisfying the following properties where $(x^{\prime},y^{\prime})=\sum_{P\in\Psi}\lambda_{P}\xi^{P}$:
  1. (i)
    $(x^{\prime},y^{\prime})\leq(x^{*},y^{*})$
  2. (ii)
    $(x^{\prime},y^{\prime})|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$
  3. (iii)
    $(\lambda,\Psi)$ minimize the quantity $\begin{array}[]{l}\\ \Delta(\lambda,\Psi):=\sum_{e\in\delta(v)}\Delta_{e}(\lambda,\Psi)\\ \text{where }\Delta_{e}(\lambda,\Psi):=y_{e}^{*}-y_{e}^{\prime}\ \forall e\in\delta(v)\end{array}$ among the ones that satisfy Conditions (i)-(ii).
Note that $(x^{*},y^{*})|_{G\setminus v}$ belong to $\mathcal{C}_{P}(G\setminus v)$, therefore the induction hypothesis implies that $(x^{*},y^{*})|_{G\setminus v}$ is obtained as an integer conic combination of induced paths of $G\setminus v$. This combination, say $(\Psi^{\prime},\lambda^{\prime})$, satisfies Conditions (i)-(ii) but is somehow the worst in terms of Condition (iii) as $\Delta(\lambda^{\prime},\Psi^{\prime})=y^{*}(\delta(v))$. Nevertheless, this proves that $(\lambda,\Psi)$ exists. For the sake of simplicity, we consider that when $\lambda_{P}=0$, $P$ does not belong to $\Psi$.
We first propose a decomposition of $\Psi$, and in the form of lemmas, we exhibit some properties satisfied by $(\lambda,\Psi)$. Given $u\in N(v)$, let us decompose $\Psi$ into the following subsets:
  • $\Psi^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and do not contain any other element of $N[v]$
  • $\Psi_{2}^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and contain another vertex $w$ in $N(v)$
  • $\Psi_{I_{2}}^{u}$ is the subset of $\Psi$ whose element’s interior contain $u$ and another element $w$ of $N(v)$.
  • $\Psi_{I_{1}}^{u}$ is the subset of $\Psi$ whose elements’s interior contain $u$ but no other vertex of $N(v)$
  • $\Psi_{u}^{v}$ is the subset of $\Psi$ whose elements contain both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
  • $\Psi_{\bot}$ contains the remaining elements.
[…]
Figure 3: Types of induced paths considered in Section 2.1
We denote by $K_{2}^{u}$ the subset of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$, that is:
$K_{2}^{u}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}.$
As the only induced path of $\Psi$ containing the edge $uv$ belong to $\Psi^{v}_{u}$ the following holds:
$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
We refer to Figure 3 for an illustration of the different types of path considered, it may be used as a visual aid to keep track of the definitions.
Lemma 2.5.
Given $u\in N(v)$, either $\Delta_{uv}(\Psi,\lambda)=0$ or $\Psi^{u}=\emptyset$.
Proof. By contradiction let $P$ be an element of $\Psi^{u}$ and $\alpha=\min\Bigl(\Delta_{uv}(\lambda,\Psi),\lambda_{P}\Bigr)$. Decreasing $\lambda_{P}$ and increasing $\lambda_{P\cup\{v\}}$ by $\alpha$ yields a new integer conic combination $(\lambda^{\prime},\Psi^{\prime})$, of induced paths’ extended incidence vectors. Moreover, as $\xi^{P}|_{G\setminus uv}=\xi^{P\cup v}|_{G\setminus uv}$, we have that $\Bigl(\sum_{P\in\Psi^{\prime}}\lambda_{P}^{\prime}\xi^{P}\Bigr)\Bigm\lvert_{G\setminus uv}=(x^{\prime},y^{\prime})|_{G\setminus uv}$. Therefore $(x^{\prime},y^{\prime})$ satisfies Conditions (i)-(ii), $\Delta_{e}(\lambda^{\prime},\Psi^{\prime})=\Delta_{e}(\lambda,\Psi)\ \forall e\in\delta(v)\setminus uv$ and $\Delta(\lambda^{\prime},\Psi^{\prime})=\Delta(\lambda,\Psi)-\alpha$. By Condition (iii) applied to $(\lambda,\Psi)$ we obtain $\alpha=0$ and as $\lambda_{P}>0$ we finally obtain $\Delta_{uv}(\Psi,\lambda)=0$. ∎
By Lemma 2.5, if $\Delta_{uv}(\Psi,\lambda)$ is nonzero, we have $\Psi^{u}=\emptyset$. Therefore, the following inequality holds and will be used in the upcoming results:
$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{u}^{v}).\\$ (9)
Lemma 2.6.
Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or $\Psi^{u}_{I_{1}}$ is nonempty.
Proof. Suppose that $\Delta_{uv}(\lambda,\Psi)\neq 0$ and that $\Psi^{u}_{I_{1}}=\emptyset$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right-hand side is equal to $\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+y_{uv}^{*}$ by definition of $(\lambda,\Psi)$. Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction. ∎
Lemma 2.7.
Given $u\in N(v)$, the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}$ are complete to $K_{2}^{u}\cup\{u\}$.
Proof. By contradiction, suppose that a neighbor $u_{1}$ of $u$ in some element $P\in\Psi^{u}_{I_{1}}$ is not adjacent to a vertex $u_{2}\in K_{2}^{u}$ contained in some induced path $P_{2}\in\Psi^{u}_{2}$. Let $\alpha=\min\Bigl(\Delta_{uv}(\lambda,\Psi),\lambda_{P},\lambda_{P_{2}}\Bigr)$. Let us consider the two inclusion wise maximal induced path contained in $P$ ending in $u$, say $_{1}, PP^{\prime}$, and let $P_{1}$ contain $u_{1}$. By Claim 2.2, $P_{1}\cup P_{2}$ induces a path of $G$. Decreasing $\lambda_{P}$, $\lambda_{P_{2}}$ and increasing $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$ of induced paths’ extended incidence vectors. Moreover, the following equality holds $\Bigl(\xi^{P}+\xi^{P_{2}}\Bigr)\Bigm\lvert_{G\setminus uv}=\Bigl(\xi^{P_{1}\cup P_{2}}+\lambda_{P^{\prime}\cup\{v\}}\Bigr)\Bigm\lvert_{G\setminus uv}$. Therefore, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(ii), $\Delta_{e}(\lambda^{\prime},\Psi^{\prime})=\Delta_{e}(\lambda,\Psi)\ \forall e\in\delta(v)\setminus\{uv\}$ and $\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})=\Delta_{uv}(\lambda,\Psi)-\alpha$. We finally get $\alpha=0$ and hence $\Delta_{uv}(\lambda,\Psi)=0$. ∎
Lemma 2.8.
Given $u\in N(v)$ either $\Delta_{uv}(\lambda,\Psi)=0$ or $\lambda_{u,v}=0$.
Proof. Suppose by contradiction that both $\lambda_{u,v}$ and $\Delta_{uv}(\lambda,\Psi)$ are nonzero. By Lemma 2.6 $\Psi_{I_{1}}^{u}$ is nonempty, let $P$ be an element of $\Psi_{I_{1}}^{u}$. $P$ contains the two distinct inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ both inducing paths ending in $u$. Let $\alpha=\min(\Delta_{uv}(\lambda,\Psi),\lambda_{P},\lambda_{\{u,v\}})$. Note that $P_{1}\cup\{v\}$ induces a path. Decreasing $\lambda_{u,v}$, $\lambda_{P}$ and increasing $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$. As $\xi^{\{u,v\}}+\xi^{P}=\xi^{\{u,v\}\cup P_{1}}+\xi^{P_{2}}$, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(ii) and $\Delta_{uv}(\lambda,\Psi)=\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})$. However, $(\Psi^{\prime})^{u}$ is nonempty as it contains $P_{2}$, by Lemma 2.5, $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction. ∎
Lemma 2.9.
Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or $K_{2}^{u}$ is nonempty.
Proof. Suppose by contradiction that $\Psi_{2}^{u}$ is empty and that $\Delta_{uv}(\lambda,\Psi)$ is strictly positive. By Lemma 2.5, Equality (9) holds. Inequality (5) associated with $u$ and the empty clique applied to $(x^{*},y^{*})$ yields: $2x_{u}^{*}\geq y^{*}(\delta(u))$. By Lemma 2.8, $\lambda_{u,v}=0$. Therefore, each $\lambda_{P}$ for $P\in\Psi_{I_{1}}^{u}\cup\Psi_{I_{2}}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$, hence the following holds:
$2x_{u}^{*}\geq 2\Bigl(\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{I_{2}}^{u})\Bigr)+\lambda(\Psi_{u}^{v})+y_{uv}^{*}.$
Equalities (8)-(9) yields $y_{uv}^{\prime}\geq y_{uv}^{*}$ implying that $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction. ∎
Lemma 2.10.
Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or the neighbors of $u$ in the elements of $\Psi_{u}^{v}$ are complete to $K_{2}^{u}\cup\{u\}$.
Proof. Suppose that $\Delta_{uv}(\lambda,\Psi)$ is nonzero. By Lemma 2.8, the elements of $\Psi_{u}^{v}$ have at least 3 vertices. Let $P=(v,u,u_{1},\dots)$ be an element of $\Psi^{v}_{u}$. By construction, $v$ is complete to $K_{2}^{u}\cup\{u\}$. Suppose that $u_{1}$ is not complete to $K_{2}^{u}\cup\{u\}$ and, without loss of generality, not adjacent to the vertex $u_{2}\in P_{2}\cap K_{2}^{u}$ for some induced path $P_{2}\in\Psi_{2}^{u}$. Let $P_{1}$ be the inclusion-wise maximal induced path included in $P$ ending in $u$ and containing $u_{1}$. By Claim 2.2, $P_{1}\cup P_{2}$ induces a path. Let $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$, decreasing $\lambda_{P}$, $\lambda_{P_{2}}$ and increasing $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$ of induced paths’ extended incidence vectors. As $\xi^{P}+\xi^{P_{2}}=\xi^{P_{1}\cup P_{2}}+\xi^{\{u,v\}}$, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(iii), and $\Delta_{uv}(\lambda,\Psi)=\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})$. Moreover, we have that $\lambda^{\prime}_{\{u,v\}}$ is nonzero, therefore by Lemma 2.8, $\Delta_{uv}(\lambda,\Psi)=0$. ∎
Theorem 2.11.
The extended incidence vectors of induced paths of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{P}(G)$.
Proof. Similarly to the proof of Theorem 1.1, the first step is to show that $\Delta_{e}(\lambda,\Psi)$ is equal to 0 for every $e\in\delta(v)$ for $(\lambda,\Psi)$ satisfying Conditions (i)-(iii). This is done through the use of an inequality (5) associated with $u$. Therefore, we need to exhibit the right clique of $N(u)$. We iteratively construct this clique that we denote by $K_{\Psi}$, and note $\Psi_{K_{\Psi}}$ the subset of $\Psi^{u}_{2}\cup\Psi_{I_{2}}^{u}$ whose elements contain a vertex of $K_{\Psi}$. Note that $\Psi_{K_{\Psi}}$ is supposed to be changing along to $K_{\Psi}$.
Start with $K_{\Psi}=K_{2}^{u}$. If there exists an element $P$ of $\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ such that the vertex $w\in\Bigl(P\cap N(u)\Bigr)\setminus N(v)$ is not complete to $K_{\Psi}$, add $P$ to $\Psi_{K_{\Psi}}$ and add the other neighbor of $u$ in $P$ to $K_{\Psi}$. This procedure is applied until the neighbors of $u$ in the elements of $\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ are complete to $K_{\Psi}\cup\{u\}$.
By Lemmas 2.7-2.10, the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}\cup\Psi_{u}^{v}$ are complete to $K_{2}^{u}\cup\{u\}$. Moreover, by the previous construction, the elements of $\Psi^{u}_{I_{2}}\setminus\Psi_{K_{\Psi}}$ satisfy this property for the whole clique $K_{\Psi}\cup\{u\}$.
It remains to show that the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}\cup\Psi_{u}^{v}$ are complete to $K_{\Psi}\setminus K_{2}^{u}$. We do so by reorganizing the elements of $\Psi_{K_{\Psi}}$.
Each vertex of $K_{\Psi}\setminus K_{2}^{u}$ is contained in an element of $\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}}$. Let $P$ be an element of $\Psi_{I_{1}}^{u}\cup\Psi^{v}_{u}$ and let $P^{\prime}$ be an element of $\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$. As $P^{\prime}$ has been added to $\Psi_{K_{\Psi}}$ by the previous construction, there exists a sequence of induced path, say $(P_{1},\dots,P_{\ell})$ such that $P_{1}=P^{\prime}$, $P_{\ell}\in\Psi_{2}^{u}$ and all the other elements belong to $\Psi_{I_{2}}^{u}$, such that the vertex $u_{i}\in(N(u)\cap P_{i})\setminus K_{\Psi}$ is nonadjacent to the vertex $v_{i+1}\in K_{\Psi}\cap P_{i+1}\ \forall i\in\{1,\dots,\ell-1\}$. For each $P_{i},\ i\in\{1,\dots,\ell-1\}$, we consider the two maximal induced path $U_{i},V_{i}$ included in $P_{i}$ and ending in $u$, with $U_{i}$ containing $u_{i}$ and $V_{i}$ containing $v_{i}$, we set $V_{\ell}=P_{\ell}$. Claim 2.2 implies that $U_{i}\cup V_{i+1}\ \forall i\in\{1,\dots,\ell-1\}$ induces a path, see Figure 4 for an illustration where the dashed line represents a nonedge. Therefore, for $\alpha=min\{\lambda_{P_{i}}:i\in\{1,\dots,\ell\}\}$ decreasing $\lambda_{P_{i}}\ \forall i\in\{1,\dots,\ell\}$, increasing $\lambda_{V_{i+1}\cup U_{i}}\forall i\in\{1,\dots,\ell\-1\}$ and $\lambda_{V_{1}},$ by $\alpha$ yields a nonnegative combination of induced paths’ extended incidence vectors $(\lambda^{\prime},\Psi^{\prime})$. Moreover $\sum_{i=1}^{\ell}\xi^{P_{i}}=\xi^{V_{1}}+\sum_{i=1}^{\ell-1}\xi^{U_{i}\cup V_{i+1}}$, implying that, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(iii) and $(\Psi^{\prime})^{u}_{2}$ contains $V_{1}$. By Lemmas 2.7 and 2.10, $P$ contains two vertices adjacent to the only vertex in $V_{1}\cap K_{\Psi}$. Note that this observation can be made for each element in $\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$ and each element in $\Psi^{u}_{I_{1}}\cup\Psi^{v}_{u}$. This implies that the neighbors of $u$ in the elements of $\Psi^{u}_{I_{1}}\cup\Psi^{v}_{u}$ have two vertices complete to $K_{\Psi}\cup\{u\}$.
As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is:
$2x_{u}^{\prime}=2x_{u}^{*}\geq y^{*}\Biggl(\delta\Bigl(u,C\left(K_{\Psi}\cup\{u\}\right)\Bigl)\Biggr)+2y^{*}\Biggl(\delta\Bigl(u,K_{\Psi} \Bigr)\Biggr)$ (10)
By the previous paragraph, each element $P$ of $\Psi_{u}^{v}\cup\Psi_{I_{1}}^{u}\cup\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ contains $u$ and its neighbors in $P$ are complete to $K_{\Psi}$. This implies that $\lambda_{P}$ contributes two times in
$y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\Biggr).$
Therefore, the following holds:
$\begin{array}[]{rl}y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\Biggr)&=y^{*}_{uv}+y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\setminus\{uv\}\Biggr)\\ &=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I_{1}}^{u})+2\lambda(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}).\end{array}$
Moreover, each element of $\Psi_{K_{\Psi}}$ intersects $K_{\Psi}$ and contains $u$, implying that it contributes one time in $y^{*}\Biggl(\delta\Bigl(u,K_{\Psi} \Bigr)\Biggr)$. Using equality (9) and inequality (10) we get:
$2\Bigl(\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{u}^{v})\Bigr)\geq y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I_{1}}^{u})+2\lambda(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}})+2\Bigl(\lambda(\Psi_{K_{\Psi}})\Bigr)$
Moreover, $\Psi_{I_{2}}^{u}=(\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}})\cup(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}})$ and $\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}}=\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$. Therefore, $y_{uv}^{\prime}\geq y_{uv}^{*}$ by equality (8). This reasoning can be made for each $uv\in\Delta(v)$. Therefore, $\Delta(\lambda,\Psi)=0$. Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired nonnegative integer combination of $(x^{*},y^{*})$ with induced paths’ extended incidence vectors. ∎
[…]
Figure 4: Structure of the paths exhibited in Theorem 2.11
Remember that $C_{4}$ is a non-chordal graph for which the extended incidence vectors of induced trees form a Hilbert basis. Therefore, as induced trees of $C_{4}$ coincide with its induced path, Theorem 2.11 is not a characterization of the graph for which the induced path extended incidence vectors form a Hilbert basis. This result also raises the question of a characterization of the graphs for which induced paths’ extended incidence vectors form Hilbert bases.
2.2 Induced paths of chordal graphs and integral systems
In this section, we discuss the consequences of Theorem 2.11 and use it to give a compact description of the induced path polytope of chordal graphs.
Corollary 2.12.
The induced path cone and polytope of chordal graphs have IDP.
Similarly to what we do for the induced tree polytope, we exhibit a description of the induced path polytope of chordal graphs. We denote by $\mathcal{P}(G)$ the polytope defined as the intersection of $\mathcal{C}_{P}(G)$ and the hyperplane:
$\{(x,y)~|~x(V)-y(E)=1\}.$ (11)
Theorem 2.13.
A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary.
Proof. $(\Rightarrow)$ The extreme points of $\mathcal{P}(G)$ are scalings of $\mathcal{C}_{P}(G)$’s generators which belong to (11). Theorem 2.11 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11). Finally, $\mathcal{P}(G)$ is included in $\mathcal{T}(G)$ which is bounded by Theorem 1.3. Therefore, $\mathcal{P}(G)$ is also bounded.
$(\Leftarrow)$ The proof is by contradiction. The same non binary point exhibited in the proof of Theorem 1.3 also belongs to $\mathcal{P}(G)$. ∎
Note that inequalities (6) are mandatory only when the graph has isolated vertices. If the input graph has no such vertex, then inequalities (5) and (7) ensure that $x$ is positive. The proof of Theorem 2.13 shows that the extreme points of $\mathcal{P}(G)$ are induced paths’ extended incidence vectors, which implies the following.
Corollary 2.14.
A graph $G$ is chordal if and only if its induced path polytope is equal to $\mathcal{P}(G)$.
It turns out that not all inequalities (5) define facets; we characterize which ones do.
Theorem 2.15.
Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\subseteq N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K\cup\{w\})]$ is bipartite.
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$ for some $l\in\{1,\dots k\}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K\cup\{w\},\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing the valid inequalities $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$. This shows that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality for a given $w$ and $K$. This is sufficient as by equality (11) the dimension of $\mathcal{P}(G)$ is at most |V|+|E|1|V|+|E|-1. We consider $K_{w}=K\cup\{w\}$. The points are as follows:
  1. 1.
    $\xi^{\{v\}}\ \forall v\in V\setminus\{w\}$
  2. 2.
    $\xi^{\{u,v\}}\ \forall uv\in E\setminus\delta(w)$
  3. 3.
    $\xi^{\{u,w\}}\ \forall uw\in\delta(w,K)$
  4. 4.
    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus\Biggl(\delta\Bigl(K\Bigr)\cup\delta\Bigl(C(K_{w})\Bigr)\Biggr)$ for some vertex $v\in K$ non adjacent to $u$
Points 1 to 4 satisfy inequality (5) for $w$ and $K$ with equality. Moreover, they can be ordered so that the corresponding matrix is triangular. For the remaining $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ elements, two edges in $\delta\Bigl(C(K_{w})\Bigr)$ are needed to achieve equality; therefore, it is not possible to keep the triangular structure.
The remaining $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$. $T_{c}$ exists and may be chosen such that $|T_{c}|=|V(\overline{G}_{c})|$: start with an odd cycle and then add the missing edges for this set to span the vertices. Now, match each edge of $\delta(w,C(K_{w}))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K_{w})|=|\delta(w,C(K_{w}))|$. Let $uw$ be in $\delta(w,C(K_{w}))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non-adjacent, hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. We refer to Figure 5 for an illustration of the matrix whose rows are the points previously exhibited and in the same order, therefore, the lines containing the submatrix $H$ correspond to the last set of $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ points. The symbol "?" stands for the submatrices associated with incidence vectors of 3-vertex paths and whose value does not need to be explicit for the validity of the proof.
$H$ is identical to the incidence matrix of the non-bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
Full row rank of the whole matrix is obtained by the facts that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and that $H$ is the maximal submatrix with no 0 row having elements of $\delta\Bigl(w,C(K_{w})\Bigr)$ as column set. This implies that the whole matrix has rank $|V|+|E|$, therefore the inequality is facet defining. ∎
[…]
Figure 5: Matrix of points built in the proof of Theorem 2.15 where $G^{\prime}$ is the graph with vertex set $V$ and edge set $(E\setminus\delta(w))\cup(\delta(w)\cap\delta(K))$
The characterization of Theorem 2.15 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently.
We introduce a necessary condition on the clique $K$ for inequality 5 to be facet-defining. After that, we show that the set of cliques verifying such a property come in a polynomial number in chordal graphs.
Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, let $K_{w}=K\cup\{w\}$, $K_{w}^{\prime}=K^{\prime}\cup\{w\}$. If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$. We say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$. Inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
Theorem 2.16.
Given a chordal graph $G$, the set of non-dominated inequalities (5) has size $O(|V|^{2})$.
Proof. We build a chordal auxiliary graph $H$ to $G[N(w)]$, whose maximal cliques are in bijection with orbit-defining cliques of $G[N(w)]$. Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will concludes. The auxiliary graph $H$ is built from $G[N(w)]$ by first adding a true twin to each vertex of $N(w)$, let $W$ be the set of added twins. Secondly, add all the edges between elements of $W$.
A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$. This proves that $H$ is chordal.
Let $K_{H}$ be a maximal clique of $H$. It is decomposed into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Conversely, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{K}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$. Summing over all $w\in V$, the total number of non-dominated inequalities (5) is at most $O(|V|^{2})$. ∎
Corollary 2.17.
There exists an extended formulation for the induced path polytope of chordal graphs with size $O(|V|^{2})$.
Corollary 2.18.
The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
The targets of Theorems 2.15 and 2.16 differ: there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 6. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of length 4 which is bipartite therefore, Theorem 2.15 implies that besides defining an orbit, its associated inequalities are not facet defining.
11223344556677
Figure 6: An orbit defining clique whose associated inequalities (5) are not facet defining
Conclusion
In this article, we show that the induced tree and path extended incidence vectors of chordal graphs respectively, form two Hilbert bases. We use this fact to give two compact linear systems of inequalities whose binaricity characterizes chordality. It turns out that there exist non-chordal graphs for which induced tree and path extended incidence vectors form a Hilbert basis, which raises the question of characterizing such graphs. These results imply that both corresponding linear optimization problems are polynomially solvable. In general, showing that some subsets of vectors of a ground set $\mathcal{V}$ forms Hilbert bases is used to show that the linear system whose rows are elements of $\mathcal{V}$ is TDI. In this work, we do not consider such systems and observe that extended incidence vectors of induced trees and path form Hilbert bases in chordal graphs. It is still a question whether these new Hilbert basis can be of any use to find new TDI systems.
Statements and Declarations
The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support and declares that he has no conflicts of interest.